# ScrollTimes: Tracing the Provenance of Paintings as a Window into History

### v1

@@ 5.2.2 Cohort View
Cohort analysis is essential for art historians to understand the relationships and interactions between historical figures [68, 69]. Matrix provides a comprehensive and efficient way to visualize these complex relationships and identify patterns. The Cohort View (Fig. 3E2-2) displays the relationship between the sealer and inscriber of a painting. We also support art historians in adding new figures to the matrix for further exploration. The type and shade of color are used to encode the event category and quantity.
5.3 Similar Painting View
Art historians hope to discover potential painting apprenticeship relationships through the similarity of the painting styles (Fig. 3E3). The approach outlined in Section 4.1.2 was utilized to obtain data on paintings that exhibit similarities to the original painting. Meanwhile, some paintings are similar to each other in terms of theme, so we use Term Frequency-Inverse Document Frequency (TF-IDF) [60] model to obtain similar painting data. In order to facilitate art historians in comparing paintings, we set the view position to be close to the original painting and sort the paintings according to the author’s birth year. When an art historian is interested in a similar painting, they can click on the [Uncaptioned image]Handscroll Feature View button to start a new exploration of the painting. Similarly, if they are interested in the author of a similar painting, they can click on either the [Uncaptioned image]Linked Data View or [Uncaptioned image]Painting Biography View button to explore that painter.
5.4 Painting Biography View
The Painting Biography View demonstrates the order of collectors and appreciators of the painting along a timeline (Fig. 3E4) (R3).
@@ 5.4 Painting Biography View
Mapping figure type with shape and color. The figures displayed in the view are divided into three types (Fig. 5). The painting’s author is represented by a circular seal. The collectors and appreciators are represented by a square seal with the same color code as in Fig. 3E1-1. Figures complemented by art historians have a dashed border.
@@
Visualizing the life span with handscroll. We use the handscroll as a metaphor for the figure’s life span (Fig. 5), which term the life-handscroll. Above the life-handscroll, we use black blocks to encode inscriptions with exact time, while the size of the blocks encodes the number of words in each inscription. Below the life-handscroll, we use red squares to encode seals with the exact time. The untimed inscriptions and seals are placed on the left side of the life-handscroll, and the number in the red square represents the number of seals. We encode common events as a line with a clickable diamond and a stacked bar chart is designed to facilitate the analysis of the distribution and types of events among figures, which can provide valuable insights into the social and cultural contexts of the figures’ lives.
@@
Figure 5: Each figure icon in the Painting Biography View encodes the figure’s basic information and its comprehensive level, and related events between figures.
@@
Displaying figure importance with hierarchy. Under the guidance of art historians, we have developed a comprehensive evaluation metric for determining the importance of figures. To evaluate a figure’s importance score $S$, we consider three key factors:
@@
The complete score of the figure is finally defined by $S=\lambda_{1}R+\lambda_{2}D+\lambda_{3}I,$ where coefficients $\lambda_{1},\lambda_{2}$, and $\lambda_{3}$ determine the weights of factors. All of the coefficients are set as 1 by default and can be adjusted by the experts according to their preference.
Interaction. We provide art historians an interactive way to integrate events into painting biographies based on their research interests (T5). Painting biographies are categorized into three trust levels: painting (analysis of the painting’s content), reasoning (based on the painting and related external information), and recommendation (automatic recommendation outside of the painting) (T6). Art historians can select the appropriate level button as per their requirements.
5.5 Uncertainty Data View
Because of the inherent characteristics of historical data, a huge number of historical figures lack comprehensive information regarding birth and death dates. To ensure the integrity of biographies, such uncertain data has been retained in the Uncertainty Data View (Fig. 3E5).
@@ 6 Evaluation
In this section, we show the usefulness and usability of ScrollTimes through two case studies and interviews with experts.
6.1 Case studies
We invited the art historians mentioned in Section 3 to freely explore ScrollTimes according to their research interests. We encouraged them to use the think-aloud protocol, and we recorded their usage of the system in two cases.
@@ 6.1 Case studies
Figure 6: Exploration Process. The scene and the figures in the painting “The Elegant Gathering in the Western Garden” have been explored (A). A1 recognizes the main figures’ information through the word cloud (A4) and the inscriptions (B), and the relationship network between the figures indicate that they have formed a close cohort (C). A1 then explores all figures’ life-scroll to study this idealized gathering (D). Through the Similar Painting View, A1 finds that “the Elegant Gathering” becomes an influential theme in future generations (E).
6.1.1 Case 1: Autumn Colors on the Que and Hua Mountains
@@
A4 has a particular fondness for the paintings of Zhao Mengfu, an outstanding painter in the Yuan Dynasty. A4 chooses “Autumn Colors on the Que and Hua Mountains”, one of the most extraordinary paintings by Zhao, to dig deep into the painting’s biography.
@@
What does the painting depict? By rotating the circularly-projected handscroll, A4 put the core painting to the top-center of vision (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryA). Consequently, there are mountains with distinct shapes (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryA1). According to the hint through the word cloud below (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryA2), it is revealed that the mountain on the left is Que Mountain, while the mountain on the right side represents Huabuzhu Mountain. It is exactly where the origin of the painting’s title comes from.
@@
How was the painting created? To gain in-depth insights into the underlying motivation for its creation, A4 subsequently analyzed the inscription left by the painter Zhao Mengfu (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryB). It recorded that Zhao Mengfu resigned from his official position in Qizhou (Jinan City, Shandong Province, China) in 1295 (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryB1). In fact, the linked external dataset CBDB also contains a recorded event corresponding to the description, which, however, lacks the annotations of starting and ending years (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryB2). After combining the two pieces of information, A4 proposed that the missing data could be filled in. It turns out that in-depth analysis based on linked data can even feed back into the complementary of the linked external databases. Zhao Mengfu’s inscription also showed that this painting was created for a figure named Gong Jin (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryB3), whose frequent appearance in the word cloud also caught A4’s interest. However, A4 was not familiar with this figure. By leveraging the Linked Data View, A4 found that Gong Jin was an alias of Zhou Mi, a celebrated poet of the Southern Song Dynasty known by other aliases Bianyang Laoren (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryB4). These terms also emerged frequently in the word cloud, which is now known as other references to Zhou Mi. Further inquiry into the historical context of the painting revealed that Zhou Mi and Zhao Mengfu were indeed close friends (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryB4).
@@
Who collected and appreciated this painting? After creation, a famous painting will be widely collected and appreciated by future generations. A4 utilized the Biography View to trace the painting’s biography from ancient times to the present (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryC).
@@
The Biography View displays the involved historical figures who left seals and inscriptions on the handscroll. As A4 noticed, Zhou Mi was absent since he did not leave any seal or inscription on the painting. However, as previously mentioned, the painting was created for Zhou Mi, which means Zhou Mi still plays a vital role in the painting’s biography. A4 thus added Zhou Mi to the Biography View manually (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryC1). The Biography View implies that this painting has been passed down through the Yuan, Ming, and Qing dynasties and has been collected or appreciated by fifteen historical figures. Accordingly, a few connections of common events exist among those figures, where the [Uncaptioned image]Political and [Uncaptioned image]Academic relations are most representative. It is worth noting that three Chinese emperors also left seals on the painting, with Emperor Qianlong leaving up to 33 seals and 9 inscriptions. A4 found something interesting: Emperor Qianlong pointed out a mistake in the painter’s inscription. Zhao Mengfu incorrectly stated the position between Que Mountain and Hua Mountain (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryC2, B5). However, after examining the content of the inscriptions left by Emperor Qianlong in detail, A4 found that they all revolved around Qianlong’s experience of visiting Qizhou and viewing the Que and Hua Mountains (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryC3). His repeated inscriptions on the same event did not provide much valuable information. Therefore, A4 chose to lower his importance (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryC4). From the above steps, A4 safely drew the collection and appreciation path of this painting, which serves as the most essential part of the painting biography. Nevertheless, A4 tried to investigate the painting genres to complement the biography.
@@
Who is influenced by this painting? The painting style and skills will be passed on, affecting other people, which is also a crucial part of the painting’s biography. Specifically, A4 found that a certain similar painting style occurs in other paintings from Huang Gongwang (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryA3), Ni Zan, Wang Hui, and Tang Yin, etc. It inspired A4 to take an in-depth look at the Linked Data View, and A4 confirmed that Huang Gongwang did have a teacher-student relationship with Zhao Mengfu. Therefore, A4 added Huang Gongwang to the biography. So far, A4 indicated that he had a sufficient and comprehensive understanding of this painting, and had constructed its corresponding biography. As A4 stated, this handscroll, which has lived through three dynasties over for 700 years, has been connected to at least 20 influential figures. This handscroll is like a time capsule that holds many records of history.
6.1.2 Case 2: Conversations span the millenniums
@@
A1 is an art historian who studies figure paintings, and he has a particular fondness for paintings with the theme of “Elegant Gathering”. To learn more about this theme, A1 used ScrollTimes to perform a thorough analysis of the masterpiece painting “The Elegant Gathering in the Western Garden” to better comprehend the perspective it portrays and the story surrounding it.
@@
A close cohort of literati. A1 rotated the painting circle and observed this painting, in which the figures waved their writing brushes and painted, recited poems and lyrics, and meditated and received Zen instruction (Fig. 6A1). A1 looked at the seals and inscriptions extracted and identified from the painting, which didn’t arouse his research interest. However, Qianlong left many inscriptions providing important information, that is, the names of the figures in the painting. A1 viewed the word cloud and the inscriptions to further explore the figures in the painting and the story behind their gathering in the Western Garden. Combined with the word cloud (Fig. 6A4), A1 recognized that the main figures in the painting include Wang Shen, Su Shi, Su Zhe, Huang Tingjian, etc. The inscriptions show detailed information about the figures in the painting, and A1 learned that the story is that Wang Shen invited a painter to make records of the incident where he and his friends gathered in the Western Garden (Fig. 6B). The figures in the painting include well-known Su Shi and Huang Tingjian, but A1 was unfamiliar with others, so he placed these figures in the Linked Data View to explore related information. The Linked Data View revealed that every figure in the painting had affairs with Su Shi. The relationship network between the figures indicated that they had formed a close cohort (Fig. 6C). Further, through event color classification, it can be seen that most of the events occurring within the cohort were [Uncaptioned image]Academic events. Among them, there were also [Uncaptioned image]Paint interactions between Li Gonglin and Li Zhiyi, Wang Shen and Huang Tingjian. A1 realized that this painting shows a close cohort of literati with Su Shi as the core. This cohort has strong academic resources, promoting literary and artistic innovation since the mid and late Northern Song Dynasty.
@@
An imaginary gathering. Analysis of the close cohort aroused A1’s curiosity, so A1 raised a question: Is it possible for so many scholars and celebrities to gather together at the same time? To explore this answer, A1 added figures to the Biography View to examine the possibility of the intersection of this cohort of literati (Fig. 6D). The Biography View shows the birth and death dates of all the figures on the painting, all of whom lived in one era. A1 analyzed, “In terms of time, the figures in the painting were given the opportunity to gather on one day.” A1 indicated that this form of the gathering was reminiscent of the masterpiece of Western art history, Raphael’s “The School of Athens”. Raphael did not dwell on the real Plato Academy in history when he painted it; instead, he painted the ideal “the School of Athens” [56]. So A1 guessed, “This form of expression may coincide with this painting, although the time allows this gathering, but whether it is a real gathering needs further historical evidence to support.”
@@
A legacy across the ages. In order to explore the impact on other paintings and the potential painting apprenticeship relationships between paintings, A1 used the Similar Painting View to observe similar paintings (Fig. 6E). Among theme similar paintings, A1 found three paintings with the same theme of “Elegant Gathering”, across the Song, Yuan, and Ming dynasties, painted by Li Gonglin, Liu Songnian and Tang Yin (Fig. 6E). A1 indicated that “Elegant Gathering” was a famous theme in the history of painting, so it was natural to find some paintings with similar themes. This showed that later generations did not simply learn and imitate this painting in terms of painting technique and style, but more importantly, they revered and admired the literati collective centered on Su Shi. People’s longing for the elegance, spirit, and style of the literati and scholars of the Song Dynasty had led to the repeated imitation and interpretation of “Elegant Gathering” by later generations of painters, and the creation is not only a manifestation of their cultural aesthetics but also the ideal world they constructed. Finally, A1 constructed a painting biography by exploring the stories of the figures and gained a comprehensive understanding of “The Elegant Gathering in the Western Garden.”

### v2

@@ 5.2.2 Exploring Cohort Social Networks
We present the relationships between historical figures through matrix visualization (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryB2) (T3, T4). Cohort analysis is essential for art historians to understand the relationships and interactions between historical figures [82, 81]. Matrix diagram provides a comprehensive and efficient way to visualize these complex relationships and identify patterns. This view (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryB2) displays the relationship among the painter, sealer and inscriber of the painting. We also support art historians in adding new historical figures to the matrix for further exploration. The types of historical events that facilitate further analysis by art historians, we have coded them separately with different colors.
5.3 Provenance-Level Component
After recognizing handscroll features and relationships among historical figures, historians proceed to understand the handscroll’s provenance using the Provenance-Level Component. It offers an automatically generated biography (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryC) (R3, T5) and supports customization based on their research interests (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryD, E) (R3, T6).
5.3.1 Creating Biographies Automatically
We merge the extracted data from the Artifact-Level and Contextual-Level (Section 4.2, Section 4.3) using temporal visualization (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryC) (R3, T5). We also apply visual encoding to different data elements to enhance the efficiency of analysis for art historians.
@@ 5.3.1 Creating Biographies Automatically
Mapping historical figure type with shape and color. Historical figures in the view are categorized into three types. The painting’s author is represented by a circular seal (Fig. 5A2). Collectors and appreciators are denoted by a square seal (Fig. 5A1) with the same color code as in ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryA1. Historical figures complemented by art historians have a dashed border (Fig. 5A3).
@@
Visualizing the historical figure’s lifespan. For the visual consistency of the entire system, we employ the visual representation of the handscroll to symbolize the historical figure’s life span, which we refer to as the “life-handscroll” (Fig. 5B1). Above the life-handscroll, we use black blocks to encode inscriptions with exact time, while the size of the blocks encodes the number of words in each inscription. Below the life-handscroll, we use red squares to encode seals with the exact time (Fig. 5B2). The untimed inscriptions and seals are placed on the left side of the life-handscroll, and the number in the red square represents the number of seals (Fig. 5B3). We encode common events as a line with a clickable diamond (Fig. 5B4) and a stacked bar chart is designed to facilitate the analysis of the distribution and types of events among historical figures, which can provide valuable insights into the social and cultural contexts of the historical figures’ lives (Fig. 5B5).
@@
Figure 5: Each historical figure icon in the Painting Biography View encodes the figure’s basic information and its comprehensive level, and related events between figures.
@@
Displaying historical figure importance with hierarchy. Under the guidance of art historians, we have developed a comprehensive evaluation metric for determining the importance of figures (Fig. 5C). To evaluate a figure’s importance score $S$, we consider three key factors:
@@
The complete score of the figure is finally defined by $S=\lambda_{1}R+\lambda_{2}D+\lambda_{3}I,$ where coefficients $\lambda_{1},\lambda_{2}$, and $\lambda_{3}$ determine the weights of factors. All of the coefficients are set as 1 by default and can be adjusted by the experts according to their preference. When figure icons overlap, we provide a hover interaction that highlights and pops them out, reducing visual clutter.
5.4 Creating Biographies Interactively
We offer various interaction methods for art historians to create biographies with research interests (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryD, E) (R3, T6).
5.4.1 Similar Painting View
Art historians hope to discover potential painting apprenticeship relationships through the similarity of the painting styles (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryD). The approach outlined in Section 4.2.2 was utilized to obtain data on paintings that exhibit similarities to the original painting. Meanwhile, some paintings are similar to each other in terms of theme, so we use Term Frequency-Inverse Document Frequency (TF-IDF) [73] model to obtain similar painting data. In order to facilitate art historians in comparing paintings, we set the view position to be close to the original painting and sort the paintings according to the author’s birth year. When an art historian is interested in a similar painting, they can click on the [Uncaptioned image]Handscroll Feature View button to start a new exploration of the painting. Additionally, if they are interested in the author of a similar painting, they can click on either the [Uncaptioned image]Linked Data View or [Uncaptioned image]Painting Biography View button to explore that painter.
5.4.2 Uncertainty Data View
During our interview with art historians, they all expressed interest in the uncertainty data in paintings. Specifically, when processing similar painting data, identifying artworks without artist names and dates (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryE1) can inspire further artistic style analysis. Furthermore, in the figure extraction process, historical figures rarely found in other databases (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryE2) piqued art historians’ curiosity. They noted that this uncertainty data holds research potential.
@@ 6 Evaluation
We invited six collaborating art historians (Section 3.1) and two additional art historians, A7 and A8, to assess the system. These two experts were not part of the system’s development. A7 is a Ph.D. student, and A8 is a professor with extensive art history experience.
We gave a tutorial on how to explore the ScrollTimes with example-based introductions [77]. (1) Introduction to our motivation and system design (25 minutes). (2) Free exploration of ScrollTimes using Google Chrome on a PC with a screen resolution of $1920\times 1080$ (45 minutes). We encouraged participants to use the think-aloud protocol to document their experiences. (3) One-on-one interviews to collect their feedback (20 minutes). We present two case studies and expert feedback to evaluate the usefulness and usability of ScrollTimes.
@@
Figure 6: Exploration Process. (A) The scene and the figures in the painting “The Elegant Gathering in the Western Garden” have been explored. A1 recognizes the main figures’ information through (A2) the word cloud and the (B) inscriptions , and (C) the relationship network between the figures indicate that they have formed a close cohort. (D) A1 then explores all figures’ life-handscroll to study this idealized gathering. (E) Through the Similar Painting View, A1 finds that “the Elegant Gathering” becomes an influential theme in future generations.
6.1 Case 1: Autumn Colors on the Que and Hua Mountains
@@
A1 has a particular fondness for the paintings of Zhao Mengfu, an outstanding painter in the Yuan Dynasty. A1 chooses “Autumn Colors on the Que and Hua Mountains,” one of the most extraordinary paintings by Zhao, to dig deep into the painting’s biography.
@@
What does the painting depict? By rotating the circularly-projected handscroll, A1 put the core painting to the top-center of vision (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryA). Consequently, there are mountains with distinct shapes (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryA6). According to the hint through the word cloud below (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryA2), it is revealed that the mountain on the left is Que Mountain, while the mountain on the right side represents Huafuzhu Mountain (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryA7) (T1, T2). It is exactly where the origin of the painting’s title comes from.
@@
How was the painting created? To gain in-depth insights into the underlying motivation for its creation, A1 subsequently analyzed the inscription left by the painter Zhao Mengfu (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryA3). It recorded that Zhao Mengfu resigned from his official position in Qizhou (Jinan City, Shandong Province, China) in 1295 (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryA8). In fact, the linked external dataset CBDB also contains a recorded event corresponding to the description, which, however, lacks the annotations of starting and ending years (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryB4). After combining the two pieces of information, A4 proposed that the missing data could be filled in (T3, T4). It turns out that in-depth analysis based on linked data can even feed back into the complementary of the linked external databases.
Zhao Mengfu’s inscription also showed that this painting was created for a figure named Gong Jin (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryA9), whose frequent appearance in the word cloud also caught A1’s interest. However, A1 was not familiar with this figure. By leveraging the Linked Data View, A1 found that Gong Jin was an alias of Zhou Mi, a celebrated poet of the Southern Song Dynasty known by other aliases Bianyang Laoren (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryB3) (T3, T4). These terms also emerged frequently in the word cloud, which is now known as other references to Zhou Mi. Further inquiry into the historical context of the painting revealed that Zhou Mi and Zhao Mengfu were indeed close friends (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryB3).
@@
Who collected and appreciated this painting? After creation, a famous painting will be widely collected and appreciated by future generations. A1 utilized the Biography View to trace the painting’s biography from ancient times to the present (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryC) (T5). This view displays the involved historical figures who left seals and inscriptions on the handscroll. As A1 noticed, Zhou Mi was absent since he did not leave any seal or inscription on the painting. However, as previously mentioned, the painting was created for Zhou Mi, which means Zhou Mi still plays a vital role in the painting’s biography. A1 thus added Zhou Mi to the Biography View manually (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryC1).
@@
The Biography View implies that this painting has been passed down through the Yuan, Ming, and Qing dynasties and has been collected or appreciated by fifteen historical figures. Accordingly, a few connections of common events exist among those figures, where the [Uncaptioned image]Political and [Uncaptioned image]Academic relations are most representative. It is worth noting that three Chinese emperors also left seals on the painting, with Emperor Qianlong leaving up to 33 seals and 9 inscriptions. A1 found something interesting: Emperor Qianlong pointed out a mistake in the painter’s inscription. Zhao Mengfu incorrectly stated the position between Que Mountain and Hua Mountain (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryC2, C4). However, after examining the content of the inscriptions left by Emperor Qianlong in detail, A1 found that they all revolved around Qianlong’s experience of visiting Qizhou and viewing the Que and Hua Mountains (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryC3). His repeated inscriptions on the same event did not provide much valuable information. Therefore, A1 chose to lower his importance (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryC5) (T6). From the above steps, A1 safely drew the collection and appreciation path of this painting, which serves as the most essential part of the painting biography. Nevertheless, A1 tried to investigate the painting genres to complement the biography.
@@
Who is influenced by this painting? The painting style and skills will be passed on, which is also a crucial part of the painting’s biography. Specifically, A1 found that a certain similar painting style occurs in other paintings from Huang Gongwang (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryD2), Ni Zan, and Tang Yin, etc (T6). It inspired A1 to take an in-depth look at the Linked Data View, and A1 confirmed that Huang Gongwang did have a teacher-student relationship with Zhao Mengfu. They were in the same painting genre. Therefore, A1 added Huang Gongwang to the biography.
So far, A1 indicated that he had a comprehensive understanding of this painting and had constructed its corresponding biography. As A1 stated, this handscroll, which has endured for over 700 years, has been connected to at least 20 influential figures. This handscroll is a time capsule holding many historical records.
6.2 Case 2: Conversations span the millenniums
@@
A5 is an art historian who studies figure paintings, and he has a particular fondness for paintings with the theme of “Elegant Gathering.” To learn more about this theme, A5 used ScrollTimes to perform a thorough analysis of the masterpiece painting “The Elegant Gathering in the Western Garden” to better comprehend the perspective it portrays and the story surrounding it.
@@
A close cohort of literati. A5 rotated the painting circle and observed this painting, in which the figures waved their writing brushes and painted, recited poems and lyrics, and meditated and received Zen instruction (Fig. 6A1) (T2). A5 looked at the seals and inscriptions extracted and identified from the painting, which didn’t arouse his research interest. However, Qianlong left many inscriptions providing important information, that is, the names of the figures in the painting. A5 viewed the word cloud and the inscriptions to further explore the figures in the painting and the story behind their gathering in the Western Garden. Combined with the word cloud (Fig. 6A4), A5 recognized that the main figures in the painting include Wang Shen, Su Shi, Su Zhe, Huang Tingjian, etc (T1, T2). The inscriptions show detailed information about the figures in the painting, and A5 learned that the story is that Wang Shen invited a painter to make records of the incident where he and his friends gathered in the Western Garden (Fig. 6B).
The figures in the painting include well-known Su Shi and Huang Tingjian, but A5 was unfamiliar with others, so he placed these figures in the Linked Data View to explore related information (T4, T5). The Linked Data View revealed that every figure in the painting had affairs with Su Shi. The relationship network between the figures indicated that they had formed a close cohort (Fig. 6C). Further, through event color classification, it can be seen that most of the events occurring within the cohort were [Uncaptioned image]Academic events. Among them, there were also [Uncaptioned image]Paint interactions between Li Gonglin and Li Zhiyi, Wang Shen and Huang Tingjian. A5 realized that this painting shows a close cohort of literati with Su Shi as the core. This cohort has strong academic resources, promoting literary and artistic innovation since the mid and late Northern Song Dynasty.
@@
An imaginary gathering. Analysis of the close cohort aroused A5’s curiosity, so A5 raised a question: Is it possible for so many scholars and celebrities to gather together at the same time? To explore this answer, A5 added figures to the Biography View to examine the possibility of the intersection of this cohort of literati (Fig. 6D) (T6). The Biography View shows the birth and death dates of all the figures on the painting, all of whom lived in one era (T5). A5 analyzed, “In terms of time, the figures in the painting were given the opportunity to gather on one day.” A5 indicated that this form of the gathering was reminiscent of the masterpiece of Western art history, Raphael’s “The School of Athens.” Raphael did not dwell on the real Plato Academy in history when he painted it; instead, he painted the ideal “the School of Athens” [66]. So A5 guessed, “This form of expression may coincide with this painting, although the time allows this gathering, but whether it is a real gathering needs further historical evidence to support.”
@@
A legacy across the ages. In order to explore the impact on other paintings and the potential painting apprenticeship relationships between paintings, A5 used the Similar Painting View to observe similar paintings (Fig. 6E) (T6). Among theme similar paintings, A5 found three paintings with the same theme of “Elegant Gathering,” across the Song, Yuan, and Ming dynasties, painted by Li Gonglin, Liu Songnian and Tang Yin (Fig. 6E). A5 indicated that “Elegant Gathering” was a famous theme in the history of painting, so it was natural to find some paintings with similar themes. This showed that later generations did not simply learn and imitate this painting in terms of painting technique and style, but more importantly, they revered and admired the literati collective centered on Su Shi. People’s longing for the elegance, spirit, and style of the literati and scholars of the Song Dynasty had led to the repeated imitation and interpretation of “Elegant Gathering” by later generations of painters, and the creation is not only a manifestation of their cultural aesthetics but also the ideal world they constructed.
Finally, A5 constructed a painting biography by exploring the stories of the figures and gained a comprehensive understanding of “The Elegant Gathering in the Western Garden.”